OpenStudy (anonymous):

Find the distance between the points given. (-1, -1) and (1, 3)

OpenStudy (shinalcantara):

the distance of two points can be determined using the formula: \[d = \sqrt{(x_2-x_1)^2 + (y_2 - y_1)^2}\] having the points, you'll have: y2 = 3 y1 = -1 x2 = 1 x1 =-1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i got 18

OpenStudy (shinalcantara):

\[d = \sqrt{(1-(-1))^2 + (3-(-1))^2}\] \[d = \sqrt{(2)^2 +(4)^2}\] \[d = \sqrt{20}\] or \[d = 2\sqrt{5}\] or in decimal \[d = 4.47 units\]